please find the value of x:
(x+2)^3/2=-1

Answer by Alan3354(69443)

(x+2)^3/2=-1
Square both sides
(x+2)^3 = 1
Find the cube roots
x+2 = 1
x = -1
============
There are 2 complex roots, also.